### v1

@@ 7.2 Potential Failures of Scaling Law
The usual experience taught that a larger data size helps neural networks break the information bottleneck to get a better excess risk bound. However, there exists a probability that including more data leads to the noise level $\xi$ growing with the dataset size, we denote it as $\xi({\sf N})$. When $\xi({\sf N})\propto{\sf C}^{1/10}$ or $\xi({\sf N})\propto{\sf N}^{1/2}$, the upper bound on the excess risk degrades to $O(1)$, therefore, such a scaling law would generally become abolished.
The Waste of Model Size Scaling. We review the boundary ${\sf M}<{\sf N}^{\frac{3}{2\zeta}}$ in Theorem 7.2. This suggests that the Model-Law (improving performance by adding parameters) is only effective when the model is small relative to the dataset’s latent complexity. Once the model becomes excessively large (in the highly over-parameterized regime), the risk is limited by the dataset size $\sf N$} rather than the parameter count $\sf M$}.
@@ Acknowledgement
Appendix A Technical Preliminary
A.1 Notations
@@
A.2 Original Model Definitions
Definition A.1 (Weights and Initialization).
@@ A.2 Original Model Definitions
Definition A.2 (Data Distribution).
@@
Definition A.3 (Original Dataset).
@@
Definition A.4 (Model Functions).
@@
A.3 Basic Facts and Lemmas
Fact A.7.
@@ A.3 Basic Facts and Lemmas
Lemma A.14 (Hoeffding bound).
@@
Lemma A.15 (Markov’s inequality).
@@
Lemma A.16 (Chernoff bound).
@@
Appendix B Gradient Computation and Learning Dynamics
B.1 Simplified Model Definitions
Definition B.1 (Rearranged Dataset).
@@ B.1 Simplified Model Definitions
Definition B.2 (Simplified Model Function).
@@
B.2 Gradient Computation
Lemma B.4.
@@ B.3 Learning Dynamics
Lemma B.6 (Formal version of Lemma 4.1).
@@
where the first step follows from the chain rule, the second step follows from the gradient flow (Eq. (2)), the third step follows from Part 1, Part 2 and Part 3 of Lemma B.4 and the definitions of $\beta_{(\nu),p}(t)$ and $\gamma_{(\nu),p}(t)$, the last step follows from the definition of $H_{(\nu)}(t)$. ∎
Appendix C Toolkit: Helpful Boundaries
@@
Lemma C.1.
Denote failure probability $\delta\in(0,0.1)$. Define $B:=\max\{O(\sqrt{\log(Lmd/\delta)}),1\}$. Assuming there exists a constant $R\in(0,1)$ satisfying $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$ for $r\in[m]$ and $\nu\in[N]$ and $R\in(0,1)$.
@@
Proof of Part 1. This proof follows from the initialization of $W_{(\nu)}(0)$ and the union bound of the tail bound of the Gaussian distribution (Fact A.7).
Proof of Part 2. This proof follows from initialization of $U_{(\nu)}(0)$ and the union bound of the tail bound of the Gaussian distribution (Fact A.7).
@@
where the first step follows from the definition of $\sigma_{(1),i}(t)$ and Fact A.12, the second step follows from $\|U_{(1)}(t)-U_{(1)}(0)\|_{F}\leq R$, $\|\Lambda_{(0),i}(0)\|_{F}=O(\sqrt{L})$, $\|\Lambda_{(0),i,\ell}(0)\|_{2}=O(1)$ and Cauchy-Schwarz inequality.
@@
where these steps follow from some basic algebra, Fact A.12 and Part 10 of this Lemma.
For a certain index $k\in[d]$, we apply Hoeffding’s inequality (Lemma A.14) to each random variable $a_{(1),r}\cdot\phi(\langle w_{(1),t}(t),o_{(1),i}(t)\rangle)$, besides, we have:
@@
where the first step follows from simple algebra and Fact A.12, the second step follows from a basic tensor trick, the third step follows from simple algebra, triangle inequality, Part 4 of this lemma and $\|\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i,\ell}(0)\|_{\infty}\leq O(R/(Bd\sqrt{Ld}))$.
@@
where these steps follow from some basic algebra, Fact A.12 and Part 10 of this Lemma.
Here, we similarly apply Hoeffding inequality (Lemma A.14), and get: with a probability at least $1-\delta$, we have:
@@
where the first step follows from Lemma B.3, the second step follows from simple algebras, the third step follows from the Cauchy-Schwartz inequality, the last two steps follow from Definition A.3 and simple algebras.
@@
Then, by Hoeffding inequality (Lemma A.14), we have:
@@
Appendix D Kernel Perturbation
@@
Lemma D.1 (Formal version of Lemma 5.3).
@@
Proof of Part 1. First, we assume the condition $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$ as Lemma C.1, which will be confirmed as a provable property in the further analysis.
@@
where the first step follows from the definition of $Q_{(\nu),p,q,1}$ and Cauchy-Schwarz inequality, the second step follows from $\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}\leq 1$, the third step follows from Cauchy-Schwarz inequality, the last step follows from Part 6 and Part 13 of Lemma C.1 and $\|o_{(\nu),q}(t)\|\leq O(1)$.
@@
the event ${\sf E}_{(\nu),p,r}$ is false, since we combining Part 1, Part 9 and Part 13 of Lemma C.1, $\|o_{(\nu),q}(t)\|\leq O(1)$ and some simple algebra.
Following Fact A.10, we have:
@@
where the first step follows from Fact A.9, the second step follows from simple algebra and $\|o_{(\nu),p}(0)\|_{2}^{2}\geq\exp(-O(dB))$.
@@
Hence, using Markov’s inequality (Lemma A.15), with a probability at least $1-\delta$, we have:
@@
Following Fact A.11 and choose
@@
It is trivial to have (Part 16 of Lemma C.1):
@@
Similarly, following Fact A.11 and combining the previous result, we have:
@@
Following Fact A.13, we have:
@@
Combining Lemma B.6, Part 15 of Lemma C.1 and Fact A.11, we can show that:
@@
Bounding Gradient Norm. Following Part 3 of Lemma C.1, we have:
@@
where the first step follows from Eq. (2) and Cauchy-Schwartz inequality, the second step follows from Eq (7) and Part 15 of Lemma C.1, the third step follows from Part 14 of Lemma C.1 and the choice of $\omega$, the last two steps follow from simple algebras and choosing
@@
Tighter Bound. We plug $R=O(\frac{1}{\sqrt{m}\omega\lambda N})$ and $N=O(1)$ into Eq. (D) to get:
@@
where $C_{1}=\omega\lambda/2$ and $C_{2}=\operatorname{poly}(L,\exp(dB),1/\delta,1/\lambda)$.
Appendix E Convergence and Approximation Bound
E.1 Complexity Analysis
Lemma E.1.
@@ E.1 Complexity Analysis
E.2 Training Convergence
Theorem E.2 (Formal version of Theorem 5.5).
@@ E.2 Training Convergence
Proof of ($n,|\mathbb{B}|$)-Fixed Convergence. Following Lemma D.1, it is easy to have:
@@
Then, following Hoeffding inequality (Lemma A.14), with a probability at least $1-\delta$, we have:
@@
where $C_{1}=\omega\lambda/2$. Combining Lemma D.1, we have:
@@
E.3 Approximation
Corollary E.3 (Formal version of Corollary 6.1, approximation bound using kernel regression).

### v2

@@ 6.2 Potential Failures of Scaling Law
The standard intuition is that larger data sizes help neural networks overcome the information bottleneck, yielding a tighter excess risk bound. However, there exists a probability that including more data leads to the noise level $\xi$ growing with the dataset size, we denote it as $\xi({\sf N})$. When $\xi({\sf N})\propto{\sf C}^{1/10}$ or $\xi({\sf N})\propto{\sf N}^{1/2}$, the upper bound on the excess risk degrades to $O(1)$, therefore, such a scaling law would generally become abolished. This happens when we have to introduce some low-quality data for the sake of data expansion.
7 Case Study of GPT-2
This section showcases the validation of our theory in previous sections, which is tested on the series of GPT-2 models [24] with the TinyStories dataset [38]. It summarizes only key experimental configurations and vital results below; full details and additional results are provided in Appendix A. Section 7.1 introduces the experimental setup of the empirical evaluation, and Section 7.2 demonstrates the corresponding results.
7.1 Experimental Setups
Datasets. To facilitate controllable scaling analysis, we utilize the TinyStories dataset [38]. This synthetic dataset comprises short stories generated by GPT-3.5 [75] and GPT-4 [25], restricted to a vocabulary accessible to a typical 3-to-4-year-old. The metric of this task is the Cross-Entropy Loss. Given the total training size of over 2M and an evaluation set of 20K, we select data sizes ${\sf N}\in\{0.25\text{M},0.5\text{M},1\text{M},2\text{M}\}$ to simulate temporal scaling behavior.
Models and Training. To investigate scaling trends regarding model size, we implement the GPT-2 series [24] across four scales: (i) Small (124M), (ii) Medium (355M), (iii) Large (774M), and (iv) XL (1.5B) for language modeling fine-tuning. It is performed using the LLaMAFactory framework [103] with the AdamW optimizer [57].
Strategies for Noisy Scaling. We introduce two noise injection strategies via random token masking. The 2M training set is partitioned into four equal subsets, $S_{1},\dots,S_{4}$. We construct four cumulative datasets of increasing size, where the $k$-th dataset is the union $\bigcup_{i=1}^{k}S_{i}$. Within each subset, tokens are masked at specific rates. We use two methods to add noise: i) selected tokens are replaced by a special mask token; ii) selected tokens are replaced by random tokens that are uniformly sampled from the vocabulary. Besides, the two distinct masking schedules used are $\{0.05,0.1,0.15,0.2\}$ and $\{0.1,0.188,0.231,0.281\}$. Notably, the noise level function $\xi({\sf N})$ of the second strategy scales with $C\sqrt{\sf N}$ (for constant $C>0$), serving to validate the theoretical claims presented in Section 6.2.
7.2 Results
Visualization of Single Scaling Laws. Figure 1 demonstrates the basic scaling laws for independent variables, confirming results aligned with classical findings [55].
General Compute Scaling Law and Pattern Transition. The left panel of Figure 2 plots cross-entropy loss (log-scaled) against FLOPs, illustrating general compute scaling laws consistent with [55]. The log-scaled view in the middle panel reveals two distinct regimes: in the upper-left region, models trained longer on less data outperform those trained shorter on more data at similar FLOPs — a regime we term Longer Training Beats Data Scaling. Conversely, in the lower-right region, the opposite holds: more data with shorter training yields better - More Data Beats Time Scaling. This leads us to distinguish between a Compute-Starved Stage and a Data-Limited Stage.
We further analyze the transition between these patterns by examining the rate of loss decrease. Two representative training trajectories are shown in the right panel of Figure 2. The top subfigure shows a Time-Dominated Pattern, where validation loss for GPT2-xl trained on 2M tokens decreases super-linearly (exponent-like). The bottom subfigure illustrates a Data-Dominated Pattern, where loss diminishes nearly linearly as data scales.
$\mathsf{N^{1/2}}$ Noise Level Disables Data Scaling. Table 1 validates the theoretical claim that scaling laws break down when noise scales as $\xi({\sf N})\propto{\sf N}^{1/2}$. Under this schedule, increasing the dataset from 0.5M to 2.0M tokens does not improve performance; the gap between noisy and clean training remains stable or widens. In one case, loss increases (1.69 to 1.71) despite more data. This confirms that the error floor is sensitive to task complexity $\xi$, and that adding low-quality data can cause the excess risk to degrade to $O(1)$, effectively nullifying the data scaling law.
Mask Strategy Dataset Size
0.5M 1M 1.5M 2.0M
clean data 1.59 +0.00 1.51 +0.00 1.46 +0.00 1.44 +0.00
{0.05, 0.1, 0.15, 0.2} + mask token 1.73 +0.14 1.66 +0.15 1.63 +0.17 1.61 +0.17
{0.1, 0.188, 0.231, 0.281} + mask token 1.88 +0.29 1.81 +0.30 1.75 +0.29 1.73 +0.29
{0.1, 0.188, 0.231, 0.281} + all token 1.80 +0.21 1.69 +0.18 1.71 +0.25 1.67 +0.23
Table 1: Best Validation Cross-Entropy Loss of GPT2-medium on various dataset sizes with different masking strategies. The red numbers represents to the difference between the corresponding loss of noisy training and the loss on the clean dataset.
@@ Acknowledgement
Appendix A Experimental Validation with Details
A.1 Experimental Setup
  • Datasets and Evaluation. To facilitate a controllable scaling analysis, we utilize the TinyStories dataset [38]. This synthetic corpus consists of short stories generated by GPT-3.5 [75] and GPT-4 [25], restricted to a vocabulary accessible to a typical 3-to-4-year-old child. We evaluate performance using the Cross-Entropy Loss. From a total training pool of over 2M samples and an evaluation set of 20K, we select training subsets of sizes $N\in\{0.25\text{M},0.5\text{M},1\text{M},2\text{M}\}$ to simulate temporal scaling behavior across expanding data horizons.
  • Model Architectures and Training. We investigate scaling trends by implementing the GPT-2 architecture [24] across four distinct configurations: (i) Small (124M), (ii) Medium (355M), (iii) Large (774M), and (iv) XL (1.5B). Fine-tuning is conducted via the LLaMAFactory framework [103] using the AdamW optimizer [57]. The training hyperparameters are standardized as follows:
    • Learning Rate: $1\times 10^{-4}$ with a cosine decay schedule.
    • Warm-up Ratio: 0.1.
    • Batch Size: 1024.
    • Context Length: 1024 tokens.
    • Compute Framework: PyTorch with Flash Attention 2 [33, 34] for optimized inference and training.
[…]
  • Noisy Scaling Strategies. We introduce controlled noise through two random token masking strategies. The 2M training set is partitioned into four disjoint subsets, $\{\mathbb{S}_{1},\mathbb{S}_{2},\mathbb{S}_{3},\mathbb{S}_{4}\}$. We construct cumulative datasets $\mathbb{D}_{k}=\bigcup_{i=1}^{k}\mathbb{S}_{i}$ for $k\in\{1,\dots,4\}$. Noise is injected using two methods: (i) replacing tokens with a specialized [MASK] token, and (ii) replacing tokens with random samples drawn uniformly from the vocabulary. We apply two distinct masking schedules: $\mathcal{M}_{1}=\{0.05,0.1,0.15,0.2\}$ and $\mathcal{M}_{2}=\{0.1,0.188,0.231,0.281\}$. Notably, the noise level $\xi({\sf N})$ in the second strategy is designed to scale with $C\sqrt{\sf N}$ (where $C>0$), specifically to empirically validate the theoretical claims formulated in Section 6.2.
A.2 Supplementary Experimental Results
Training Loss Dynamics. We provide complete training loss dynamics of four models with four dataset scales in Figure 3 and Figure 4, which show the evaluation of model performance across fixed data budgets of 0.25M, 0.5M, 1M, and 2M tokens. These visualizations demonstrate that increasing model capacity consistently leads to faster convergence and lower terminal cross-entropy loss, confirming that larger architectures more efficiently extract patterns from a given quantity of information.
Figure 3: Training loss dynamics varied by the dataset size.
Figure 4: Training loss dynamics varied by the model size.
Complete Validation Loss Dynamics. We provide complete validation loss dynamics of four models with four dataset scales in Figure 5 and Figure 6, which isolate the impact of data volume on specific model architectures. By facetting the results by model size, these charts reveal how increasing the training set size directly shifts the loss floor downward. This behavior provides clear evidence of the data-limited regime, where the model’s performance is constrained not by its own parameter count, but by the diversity and volume of the underlying corpus. Together, these four files map the interplay between architecture size and data availability.
Figure 5: Validation loss dynamics varied by the dataset size.
Figure 6: Validation loss dynamics varied by the model size.
Appendix B Technical Preliminary
B.1 Notations
@@
B.2 Original Model Definitions
Definition B.1 (Weights and Initialization).
@@ B.2 Original Model Definitions
Definition B.2 (Data Distribution).
@@
Definition B.3 (Original Dataset).
@@
Definition B.4 (Model Functions).
@@
B.3 Basic Facts and Lemmas
Fact B.7.
@@ B.3 Basic Facts and Lemmas
Lemma B.14 (Hoeffding bound).
@@
Lemma B.15 (Markov’s inequality).
@@
Lemma B.16 (Chernoff bound).
@@
Appendix C Gradient Computation and Learning Dynamics
C.1 Simplified Model Definitions
Definition C.1 (Rearranged Dataset).
@@ C.1 Simplified Model Definitions
Definition C.2 (Simplified Model Function).
@@
C.2 Gradient Computation
Lemma C.4.
@@ C.3 Learning Dynamics
Lemma C.6 (Formal version of Lemma 4.1).
@@
where the first step follows from the chain rule, the second step follows from the gradient flow (Eq. (2)), the third step follows from Part 1, Part 2 and Part 3 of Lemma C.4 and the definitions of $\beta_{(\nu),p}(t)$ and $\gamma_{(\nu),p}(t)$, the last step follows from the definition of $H_{(\nu)}(t)$. ∎
Appendix D Toolkit: Helpful Boundaries
@@
Lemma D.1.
Denote failure probability $\delta\in(0,0.1)$. Define $B:=\max\{O(\sqrt{\log(Nmd/\delta)}),1\}$. Assuming there exists a constant $R\in(0,1)$ satisfying $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$ for $r\in[m]$ and $\nu\in[N]$ and $R\in(0,1)$.
@@
Proof of Part 1. This proof follows from the initialization of $W_{(\nu)}(0)$ and the union bound of the tail bound of the Gaussian distribution (Fact B.7).
Proof of Part 2. This proof follows from initialization of $U_{(\nu)}(0)$ and the union bound of the tail bound of the Gaussian distribution (Fact B.7).
@@
where the first step follows from the definition of $\sigma_{(1),i}(t)$ and Fact B.12, the second step follows from $\|U_{(1)}(t)-U_{(1)}(0)\|_{F}\leq R$, $\|\Lambda_{(0),i}(0)\|_{F}=O(\sqrt{L})$, $\|\Lambda_{(0),i,\ell}(0)\|_{2}=O(1)$ and Cauchy-Schwarz inequality.
@@
where these steps follow from some basic algebra, Fact B.12 and Part 10 of this Lemma.
For a certain index $k\in[d]$, we apply Hoeffding’s inequality (Lemma B.14) to each random variable $a_{(1),r}\cdot\phi(\langle w_{(1),t}(t),o_{(1),i}(t)\rangle)$, besides, we have:
@@
where the first step follows from simple algebra and Fact B.12, the second step follows from a basic tensor trick, the third step follows from simple algebra, triangle inequality, Part 4 of this lemma and $\|\Lambda_{(1),i,\ell}(t)-\Lambda_{(1),i,\ell}(0)\|_{\infty}\leq O(R/(Bd\sqrt{Ld}))$.
@@
where these steps follow from some basic algebra, Fact B.12 and Part 10 of this Lemma.
Here, we similarly apply Hoeffding inequality (Lemma B.14), and get: with a probability at least $1-\delta$, we have:
@@
where the first step follows from Lemma C.3, the second step follows from simple algebras, the third step follows from the Cauchy-Schwartz inequality, the last two steps follow from Definition B.3 and simple algebras.
@@
Then, by Hoeffding inequality (Lemma B.14), we have:
@@
Appendix E Kernel Perturbation
@@
Lemma E.1 (Formal version of Lemma 5.3).
@@
Proof of Part 1. First, we assume the condition $\|w_{(\nu),r}(t)-w_{(\nu),r}(0)\|_{2}\leq R$ and $\|U_{(\nu)}(t)-U_{(\nu)}(0)\|_{F}\leq R$ as Lemma D.1, which will be confirmed as a provable property in the further analysis.
@@
where the first step follows from the definition of $Q_{(\nu),p,q,1}$ and Cauchy-Schwarz inequality, the second step follows from $\operatorname{\mathbb{I}}\{\langle o_{(\nu),p}(t),w_{(\nu),r}(t)\rangle>0,\langle o_{(\nu),q}(t),w_{(\nu),r}(t)\rangle>0\}\leq 1$, the third step follows from Cauchy-Schwarz inequality, the last step follows from Part 6 and Part 13 of Lemma D.1 and $\|o_{(\nu),q}(t)\|\leq O(1)$.
@@
the event ${\sf E}_{(\nu),p,r}$ is false, since we combining Part 1, Part 9 and Part 13 of Lemma D.1, $\|o_{(\nu),q}(t)\|\leq O(1)$ and some simple algebra.
Following Fact B.10, we have:
@@
where the first step follows from Fact B.9, the second step follows from simple algebra and $\|o_{(\nu),p}(0)\|_{2}^{2}\geq\exp(-O(dB))$.
@@
Hence, using Markov’s inequality (Lemma B.15), with a probability at least $1-\delta$, we have:
@@
Following Fact B.11 and choose
@@
It is trivial to have (Part 16 of Lemma D.1):
@@
Similarly, following Fact B.11 and combining the previous result, we have:
@@
Following Fact B.13, we have:
@@
Combining Lemma C.6, Part 15 of Lemma D.1 and Fact B.11, we can show that:
@@
Bounding Gradient Norm. Following Part 3 of Lemma D.1, we have:
@@
where the first step follows from Eq. (2) and Cauchy-Schwartz inequality, the second step follows from Eq (7) and Part 15 of Lemma D.1, the third step follows from Part 14 of Lemma D.1 and the choice of $\omega$, the last two steps follow from simple algebras and choosing
@@
Tighter Bound. We plug $R=O(\frac{1}{\sqrt{m}\omega\lambda N})$ and $N=O(1)$ into Eq. (E) to get:
@@
where $C_{1}=\omega\lambda/2$ and $C_{2}=\operatorname{poly}(L,\exp(dB),1/\delta,1/\lambda)$. ∎
Appendix F Convergence and Approximation Bound
F.1 Complexity Analysis
Lemma F.1.
@@ F.1 Complexity Analysis
F.2 Training Convergence
Theorem F.2 (Formal version of Theorem 5.5).
@@ F.2 Training Convergence
Proof of ($n,|\mathbb{B}|$)-Fixed Convergence. Following Lemma E.1, it is easy to have:
@@
Then, following Hoeffding inequality (Lemma B.14), with a probability at least $1-\delta$, we have:
@@
where $C_{1}=\omega\lambda/2$. Combining Lemma E.1, we have:
@@
F.3 Approximation
Corollary F.3 (Formal version of Corollary 5.6, approximation bound using kernel regression).